### v1

# When, Why and How Much? Adaptive Learning Rate Scheduling by Refinement

Abstract
Learning rate schedules used in practice bear little resemblance to those recommended by theory. We close much of this theory/practice gap, and as a consequence are able to derive new problem-adaptive learning rate schedules. Our key technical contribution is a refined analysis of learning rate schedules for a wide class of optimization algorithms (including SGD). In contrast to most prior works that study the convergence of the average iterate, we study the last iterate, which is what most people use in practice. When considering only worst-case analysis, our theory predicts that the best choice is the “linear decay” schedule: a popular choice in practice that sets the stepsize proportionally to $1-t/T$, where $t$ is the current iteration and $T$ is the total number of steps. To go beyond this worst-case analysis, we use the observed gradient norms to derive schedules refined for any particular task. These refined schedules exhibit learning rate warm-up and rapid learning rate annealing near the end of training. Ours is the first systematic approach to automatically yield both of these properties. We perform the most comprehensive evaluation of learning rate schedules to date, evaluating across 10 diverse deep learning problems, a series of LLMs, and a suite of logistic regression problems. We validate that overall, the linear-decay schedule matches or outperforms all commonly used default schedules including cosine annealing, and that our schedule refinement method gives further improvements.
1 Introduction
For minimizing a function $f$, Stochastic Gradient Descent (SGD) updates the iterate $x_{t}$ at step $t$ via:
$x_{t+1}=x_{t}-\eta_{t}g_{t},$
where $g_{t}$ is a (possibly stochastic) sub-gradient at $x_{t}$, and $\eta_{t}$ is the learning rate (LR) at time $t$. Choosing a sequence of $\eta_{t}$ for steps $t=1,\dots,T$ is a core problem in optimization.
The learning rate sequence for an optimizer is typically decomposed into two parts: the baseline learning rate, indicating the maximum LR to use, and a schedule, a sequence that multiplies the baseline LR to give the LR sequence. In this work we focus exclusively on the problem of scheduling. Choosing the right learning rate schedule for best performance is difficult; standard practice is to perform a hyper-parameter sweep over a set of standardized schedules (Wu et al., 2020).
Setting $\eta_{t}$ from theory is difficult due to a multitude of potential problem assumptions and the wildly varying schedules that arise from these assumptions. For instance, $\eta_{t}\propto 1/\sqrt{t}$, $\eta_{t}\propto 1/t$ and constant schedules $\eta_{t}=\eta$ are all dictated by three common but different sets of assumptions. Unfortunately, all three work suboptimally in practice for deep learning (Section 3.1), and are unpopular in the community (Ge et al., 2019a). In this work we focus on two causes for this theory-practice gap:
  1. 1.
    Most theory analyzes the average iterate (Polyak, 1990; Ruppert, 1988) $\hat{x}_{T}=\frac{1}{T}\sum_{t=1}^{T}x_{t}$ or a randomly sampled iterate. However, in practice the last iterate $x_{T}$ is used.
  2. 2.
    Existing theory for the last iterate often uses crude constant bounds on the gradient norms or curvature. Our new tighter bound involves the entire gradient norm sequence instead, allowing for problem-adaptive LR schedules.
1:Input: $G=\left\|g_{t}\right\|$ sequence of length $T$, smoothing hyper-parameter $\tau>0$
[…]
3:Define $w_{t}=\hat{G}_{t}^{-2}$
4:for $t=1$ to $T$ do
5:
$\eta_{t}=w_{t}\sum_{p=t+1}^{T}w_{p}$
6:end for
7:Return normalized schedule $\eta/\max(\eta)$
Algorithm 1 Schedule Refinement for SGD
Our method is a “refinement” method: it uses a prior training run to produce an improved schedule to use in future runs. The practical variant of our schedule refinement method for SGD is given in Algorithm 1. Given a sequence of gradient norms produced by a prior run, it outputs a new schedule that is adaptive to the structure of the problem. Mathematically, it is minimizing a novel bound we derive on the function value $f(x_{T})$ of the final iterate (Section 2.1). The learning rate at each time-step involves a sum of inverse-squared gradient norms from future time-steps, a major departure from previous approaches to scheduling.
This approach is partially motivated by the remarkable work of Pan et al. (2022), who show that for the last iterate of quadratic problems, minimizing an upper bound can yield improved schedules. However their work requires knowledge of the full Eigenspectrum of the Hessian, making it impractical to use. Our theory relies on knowledge of the sequence of expected gradient norms, a more tractable quantity.
Our analytical approach also departs from previous approaches by generalizing beyond SGD. Prior analyses of learning rate schedules often rely on the particular form of the SGD iterates to drive the calculations (Jain et al., 2019; Zamani and Glineur, 2023). Our approach is instead a broad technique that provides learning rate schedules for any base optimization algorithm. On a technical level, we design a step-size schedule that converts any algorithm which obtains a vanishing regret into one that ensures a last-iterate guarantee. This means that our refinement technique can provide theoretically-motivated schedules for popular base optimization algorithms like Adam.
Overall, we make three main contributions:
  1. 1.
    Theory: We provide an general analysis of learning rate schedules for arbitrary optimization algorithms. Our approach recovers the optimal convergence rates for SGD, and can be used to produce refined schedules customized for a particular task.
  2. 2.
    Practice: Among non-adaptive schedules, we show that warm-up followed by linear decay almost always matches or outperforms other schedules, including cosine decay. Our refined schedules typically provide further improvements, suggesting that our theory provides actionable guidance even for training non-convex neural networks.
  3. 3.
    Theory meets Practice: Our refined schedules exhibit both warmup and nearly-linear decay (Figure 1). While Zamani and Glineur (2023) have very recently shown that linear decay has favorable last-iterate guarantees, to our knowledge this is the first time that warmup has arisen directly from theory rather than as an empirical heuristic (Goya et al., 2017). That is, given hindsight knowledge of all gradient statistics, our theoretically optimized schedules usually include a warmup phase during which the learning rate increases.
Figure 1: Example gradient norm sequences (top row) and the resulting refined schedules given by Algorithm 1 (bottom row). Black dashed line at $y=1$ shown for reference. Percentage of runtime on the x-axis.
1.1 Notation
$f\colon\mathbb{R}^{d}\to\mathbb{R}$ is a convex objective. $x_{1},\dots,x_{T}$ and $z_{1},\dots,z_{T}$ are random vectors in $\mathbb{R}^{d}$ with $x_{1}=z_{1}$, and $\Delta_{t}\triangleq z_{t+1}-z_{t}$. $\Delta_{t}$ will indicate “baseline” updates before applying a schedule, and $x_{t}$ will be iterates after applying the schedule. $g_{1},\dots,g_{T}$ are random vectors in $\mathbb{R}^{d}$ satisfying $\mathbb{E}[g_{t}|x_{1},\dots,x_{t}]\in\partial f(x_{t})$ ($\mathbb{E}[g_{t}]=\nabla f(x_{t})$ when $f$ is differentiable). $G^{2}$ is a bound on $\mathbb{E}[\max_{t}\|g_{t}\|^{2}]$. $w_{1},\dots,w_{T}$ indicate non-negative random variables in $\mathbb{R}$ such that $g_{t}$ and $w_{t}$ are independent given $x_{1},\dots,x_{t}$. We define $w_{a:b}\triangleq\sum_{t=a}^{b}w_{t}$. If $a>b$, then $w_{a:b}\triangleq 0$. $u$ denotes an arbitrary element of $\mathbb{R}^{d}$; typically one should consider the case $u\in\operatorname*{arg\,min}f$. $D\triangleq\|x_{1}-u\|$ is the “distance-to-solution” term, and $f_{\star}\triangleq\inf_{u}f(u)$.
2 Main Analytical Result
Our general result is Theorem 1. This allows us to convert any sequence $z_{1},\dots,z_{T}$ with bounded regret $\sum_{t=1}^{T}\langle g_{t},z_{t}-u\rangle$ into a sequence of iterates $x_{1},\dots,x_{T}$ with a bound on $f(x_{T})-f(u)$. Thus, Theorem 1 can be viewed as another addition to the family of reductions from stochastic optimization to regret bounds (Cesa-Bianchi et al., 2004; Cutkosky, 2019). The proof can be found in Appendix B.
Theorem 1
Suppose $z_{1},\dots,z_{T}$ is some arbitrary sequence of vectors. Let $w_{1},\dots,w_{T}$ be an arbitrary sequence of non-negative numbers. Recall that we define $\Delta_{t}=z_{t+1}-z_{t}$ and $x_{1}=z_{1}$. For $t\geq 1$, suppose $x_{t+1}$ satisfies:
$x_{t+1}=x_{t}+\frac{w_{t+1:T}}{w_{1:T}}\Delta_{t},$
then for any $u$:
$\mathbb{E}[f(x_{T})-f(u)]\leq\mathbb{E}\left[\sum_{t=1}^{T}\frac{1}{w_{1:T}}\langle w_{t}\cdot g_{t},z_{t}-u\rangle\right].$
Let us take a moment to consider the implications of this Theorem in the simplest setting of $w_{t}=1$ for all $t$. In this case, it is well-known that by setting $\Delta_{t}=-\eta g_{t}$ for $\eta=\frac{D}{G\sqrt{T}}$, one guarantees $\sum_{t=1}^{T}\langle g_{t},z_{t}-u\rangle\leq DG\sqrt{T}$. Thus, we immediately obtain the following important corollary:
Corollary 2
Set $x_{t+1}=x_{t}-\eta_{t}g_{t}$ with $\eta_{t}=\frac{D}{G\sqrt{T}}\left(1-\frac{t}{T}\right)$. Then:
$\mathbb{E}[f(x_{T})-f(u)]$ $\leq\frac{DG}{\sqrt{T}}.$
Proof  With $w_{t}=1$ for all $t$ and $\Delta_{t}=-\frac{D}{G\sqrt{T}}g_{t}$, classical online gradient descent analysis (Zinkevich (2003)) yields $\sum_{t=1}^{T}\langle w_{t}g_{t},z_{t}-u\rangle\leq DG\sqrt{T}$. The result now follows from Theorem 1.
The sequence $x_{t}$ suggested by Corollary 2 is simply stochastic gradient descent ($x_{t+1}=x_{t}-\eta_{t}g_{t}$) equipped with a linear decay learning rate schedule:
$\eta_{t}=\frac{D}{G\sqrt{T}}\left(1-\frac{t}{T}\right).$ (1)
Linear decay emulates the effects of iterate averaging, as the contribution from each gradient to the returned point is approximately the same as it would be in an average: the gradient $g_{T/2}$ appears in half the points in the average, and so its weight is halved, the gradient $g_{t}$ appears in $T-t$ out of $T$ points and so its weight is $1-t/T$.
This bound has a significantly better constant than previous schedules for last-iterate convergence of SGD in this setting (Jain et al., 2019), and matches recent work by Zamani and Glineur (2023), who were the first to show that this schedule is actually optimal for gradient descent for the Convex $G$-Lipschitz complexity class. Our regret analysis recovers their result when specialized to SGD.
In practice, the linear decay schedule is employed not only with SGD but also with a diverse panoply of other optimization algorithms. Our Theorem 1 suggests a theoretical basis for this approach: so long as the underlying optimization method ensures a regret bound, the linear decay schedule will provide a last-iterate guarantee. Note that we do not require the regret bound to be analytically proven (as is the case for e.g.  SGD (Zinkevich, 2003), AdaGrad (Duchi et al., 2011; Streeter and McMahan, 2010) or AMSGrad (Reddi et al., 2018)); it suffices for the regret bound to hold in practice (as may hold for Adam (Kingma and Ba, 2015) or AdamW (Loshchilov and Hutter, 2017a)).
The key initial step in the proof of Theorem 1 is an identity we call the “all-tail summation bound”. This is a refinement of a bound from Orabona (2020) who attributes the approach originally to Lin et al. (2016). Using the notation $q_{t}=f(x_{t})-f_{*}$, Lin et al. (2016) use the bound:
$w_{T}q_{T}\leq\frac{1}{T}\sum_{t=1}^{T}w_{t}q_{t}+\sum_{k=1}^{T-1}\frac{1}{k(k+1)}\sum_{t=k+1}^{T}w_{t}\left(q_{t}-q_{k}\right),$
whereas our refined approach uses an equality:
$q_{T}=\frac{1}{w_{1:T}}\sum_{t=1}^{T}w_{t}q_{t}+\sum_{k=1}^{T-1}\frac{\eta_{k}}{\sum_{t=k+1}^{T}\eta_{t}}\left(\frac{1}{\sum_{t=k}^{T}w_{t}}\sum_{t=k}^{T}w_{t}\left(q_{t}-q_{k}\right)\right).$
This identity is derived in Appendix A. All our results arise from careful manipulation of this identity.
2.1 Optimizing the bound for data-dependent schedules
We have now seen that setting $w_{t}=1$ for all $t$ recovers the linear decay schedule, and can obtain the worst-case optimal convergence rates. However, optimizing for the worst case usually yields overly pessimistic behavior. In this section, we build more adaptive schedules that obtain better results on real data. To do this, we simply choose $w_{t}$ so as to optimize the bound in Theorem 1.
We focus on the particular case of SGD by setting $x_{t+1}=x_{t}-\eta_{t}g_{t}$ with $\eta_{t}=\frac{w_{t}w_{t+1:T}}{w_{1:T}}$. In the notation of Theorem 1, this corresponds to $\Delta_{t}=-w_{t}g_{t}$. For this case, we have the following result, with proof in Appendix C.
Theorem 3
Suppose that $x_{t+1}=x_{t}-\eta_{t}g_{t}$ with $\eta_{t}=\frac{w_{t}w_{t+1:T}}{w_{1:T}}$. Then we have:
$\mathbb{E}[f(x_{T})-f(u)]$ $\leq\mathbb{E}\left[\frac{1}{2\cdot w_{1:T}}\left(D^{2}+\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2}\right)\right].$ (2)
Moreover, for a fixed sequence $\|g_{1}\|^{2},\dots,\|g_{T}\|^{2}$, the value of $\frac{1}{2\cdot w_{1:T}}(D^{2}+\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2})$ is minimized by setting:
$w_{t}=\|g_{t}\|^{-2}\frac{D}{\sqrt{\sum_{p=1}^{T}\|g_{p}\|^{-2}}}.$
Theorem 3 suggests that if we knew the sequence of gradient norms ahead of time, then we could optimize the weights $w_{t}$ (and therefore the learning rate schedule $\eta_{t}$) by setting $w_{t}\propto\|g_{t}\|^{-2}$. This yields a simple practical approach for refining a learning rate schedule based on empirical observations. First, perform one run using a baseline schedule to observe the sequence of gradient norms. Then, use these norms to compute an optimal schedule via Theorem 3. The constant factor $D=\|x_{1}-u\|$ appearing in the value for $w_{t}$ plays the role of the “scale factor” typically applied to learning rate schedules. A line of recent work has shown that this quantity can be efficiently estimated online without significant downsides (Mcmahan and Streeter, 2012; Orabona and Pál, 2016; Cutkosky and Orabona, 2018; Mhammedi and Koolen, 2020; Zhang et al., 2022; Carmon and Hinder, 2022; Ivgi et al., 2023; Khaled et al., 2023; Cutkosky et al., 2023).
There are some subtleties here. Allowing $w_{t}$ to depend on random variable $g_{t}$ breaks the independence between $w_{t}$ and $g_{t}$. Further, the act of changing the schedule from a baseline to a refined data-dependent schedule will change the gradient norm sequence, which may in turn indicate that a substantially different schedule would have been optimal. Our approach relies on the practical assumption that the gradient norm sequence does not change significantly after refinement. To encourage this, Algorithm 1 applies a median smoothing filter to the gradient norms before refining.
Finally, Theorem 3 provides an analysis that recovers schedules for SGD. However, practitioners commonly use optimization methods with more complicated per-coordinate updates or other preconditioning schemes such as Adam. In Appendix E we provide an alternative version of Theorem 3 that applies to such algorithms (Theorem 8). This result enables us to develop schedules tailored to any optimization algorithm. However, actually building this schedule in practice may require inspecting the algorithm’s internal state, which may be difficult or inefficient. For per-coordinate algorithms like Adam, we suggest simply setting $w_{t}\propto 1/\|g_{t}\|_{1}$ as an approximation (see Section 3).
Figure 2 gives the Refined schedules on a set of standard benchmark machine learning problems when initially trained using linear decay schedules to provide the gradient norm sequences. Full details of the models and training setup for each problem is provided in the Appendix. Gradient $\ell_{2}$ norms are shown for SGD trained problems (ImageNet, RCNN) and $\ell_{1}$ norms for the Adam trained problems, which also use inverse-$\ell_{1}$ norm weights for refinement (see Section 3). Our single hyper-parameter $\tau=0.1$ was tuned so that the resulting schedules obtained a balance between smoothness and capturing structure. We recommend setting this parameter by eye rather than by grid search.
Figure 2: Gradient Norm sequences and the resulting Refined schedules, generated using an initial linear decay schedule with warmup for the initial run. Log scale views are inset for scale reference. Fitted polynomial decay schedules are overlayed in blue.
3 Experiments
Table 1: Logistic Regression Experiments (Train Error Rate %).
Problem Stepwise Cosine Linear Refined $\left\|g\right\|_{2}^{2}$ Refined $\left\|g\right\|_{1}$ Refined
Aloi $13.38$$\pm 0.05$ $12.69$$\pm 0.05$ $12.76$$\pm 0.06$ $\mathbf{11.75}$$\pm 0.01$ $12.24$$\pm 0.04$ $12.30$$\pm 0.05$
Glass $31.39$$\pm 0.16$ $30.82$$\pm 0.22$ $30.72$$\pm 0.18$ $\mathbf{30.05}$$\pm 0.44$ $\mathbf{29.62}$$\pm 0.37$ $\mathbf{30}$$\pm 0.41$
Iris $\mathbf{1.39}$$\pm 0.000$ $\mathbf{1.39}$$\pm 0.000$ $\mathbf{1.39}$$\pm 0.000$ $\mathbf{1.46}$$\pm 0.07$ $\mathbf{1.46}$$\pm 0.07$ $\mathbf{1.46}$$\pm 0.07$
Letter $22.24$$\pm 0.008$ $\mathbf{22.24}$$\pm 0.01$ $\mathbf{22.20}$$\pm 0.02$ $\mathbf{22.23}$$\pm 0.03$ $\mathbf{22.20}$$\pm 0.03$ $\mathbf{22.20}$$\pm 0.03$
Pendigits $4.70$$\pm 0.02$ $4.67$$\pm 0.01$ $\mathbf{4.62}$$\pm 0.03$ $\mathbf{4.56}$$\pm 0.02$ $\mathbf{4.58}$$\pm 0.02$ $\mathbf{4.56}$$\pm 0.04$
Sensorless $11.84$$\pm 0.09$ $11.30$$\pm 0.09$ $11.29$$\pm 0.09$ $10.71$$\pm 0.08$ $\mathbf{10.08}$$\pm 0.05$ $\mathbf{10.11}$$\pm 0.05$
Vehicle $18.83$$\pm 0.09$ $18.49$$\pm 0.05$ $18.55$$\pm 0.06$ $\mathbf{18.21}$$\pm 0.12$ $\mathbf{18.19}$$\pm 0.08$ $\mathbf{18.21}$$\pm 0.1$
Vowel $23.43$$\pm 0.08$ $22.99$$\pm 0.09$ $22.94$$\pm 0.10$ $\mathbf{22.48}$$\pm 0.12$ $\mathbf{22.44}$$\pm 0.08$ $\mathbf{22.41}$$\pm 0.08$
To validate the effectiveness of our method on convex problems, we performed a comparison across 8 commonly used classification problems from the LIBSVM repository, with separable problems excluded (see Section 3.5). We used a logistic regression loss together with the Adam optimizer with $\beta=(0.9,0.95)$, with batch size 16 and trained for 100 epochs. The Refined schedules were calculated by using the linear decay schedule to generate the initial gradient norm sequence. Table 1 demonstrates that both the linear decay schedule and our refinement schedule consistently either match or out-perform the cosine schedule. The linear decay schedule matches the cosine schedule on every problem (up to statistical significance), and out-performs it on two problems. The Refined schedule further out-performs the linear decay schedule, either matching of exceeding it across the board.
Rather than using the Adam specific weighting for deriving the Refined schedules, it is often convenient to use other norms, particularly if existing logs are available using these other norms. In Table 1 we present results using both the $\ell_{1}$ norm and $\ell_{2}$ norm-squared weighting. Both are shown to maintain the advantages of the refinement technique, out-performing the non-refined schedules consistently. Given the ease of logging $\ell_{1}$ norms compared to the internal Adam state, we advocate for weights to be the inverse of the $\ell_{1}$ norm (no squaring) when using Adam.
3.1 Deep Learning experiments
Figure 3: Training curves on CIFAR-10 for a sweep over inverse-sqrt and inverse-linear hyper-parameters. A linear-decay schedule baseline is shown in orange. All combinations are out-performed by the linear-decay schedule.
Classical any-time learning rates schedules such as $1/\sqrt{t}$ and $1/t$ are commonly used in theoretical analysis of optimization methods, yet they are rarely used by practitioners. To illustrate why, in Figure 3, we sweep over the learning rate $\eta$ and offset $\beta$ for schedules of the form:
$\eta\frac{\beta}{\beta+t}\quad\text{and}\quad\eta\frac{\sqrt{\beta}}{\sqrt{\beta+t}}.$
A 5% duration learning rate warmup was also included in each case. Even with 60 hyper-parameter combinations tried for each family, neither are able to match the performance of the linear-decay schedule. For our deep learning experiments we implemented two commonly used practical variants of these schedules, since sweeping the $\beta$ hyper-parameter is computationally impractical. The vanilla version uses $\beta=1$, and the offset version uses $\beta$ set to match the duration of the warmup period (a common setting due to its implementation in FairSeq).
3.2 Large-scale schedule benchmark
We performed a large-scale comparison of schedules across common deep learning benchmarks:
CIFAR10
For a small-scale image-classification problem, we chose CIFAR10 (Krizhevsky, 2009). A high-performance Wide ResNet architecture was used (Zagoruyko and Komodakis, 2016). Test error percentage is reported.
CIFAR100
A more realistic yet still small scale benchmark. We used a compact DenseNet architecture (Huang et al., 2017) to increase the diversity of model architectures tested. Test error percentage is reported.
ImageNet
We used a ResNet-50 architecture (He et al., 2016) for the ImageNet (Russakovsky et al., 2015) classification task. Note that we used a standard (non-SOTA) data-augmentation pipeline following other recent comparisons in the optimization literature. Test error percentage is reported.
IWSLT14
A small-scale German-English translation task (Cettolo et al., 2014) using a LSTM model (Wiseman and Rush, 2016). Test Perplexity is reported.
GPT
A modern small (162M parameter) GPT-style auto-regressive transformer model (Radford et al., 2019) trained on the large Book-Wiki corpus (Zhu et al., 2015). Test Perplexity is reported.
RoBERTa
A masked autoencoder variant (Liu et al., 2019) also trained on the Book-Wiki corpus. Test Perplexity is reported.
ViT
A high-performance ImageNet classifier using a Vision Transformer (Dosovitskiy et al., 2021). In contrast to our ResNet-50 experiments, here we use modern data-augmentation, following standard practice for ViT training. Test error percentage is reported.
DLRM
A modern recommendation system engine (Naumov et al., 2019) trained on the open Criteo Kaggle Display Advertising dataset. Test Accuracy is reported.
MRI
A large stacked U-Net architecture (Sriram et al., 2020) trained on the fastMRI dataset (Zbontar et al., 2018), an image-to-image regression task from the medical imaging domain. Test set metric $100\cdot(1-\text{SSIM})$ is reported.
RCNN
[…]
For each problem, we first performed a sweep of learning rates on a grid [$10^{i}$, $2\times 10^{i}$, $5\times 10^{i}$] for varying $i$, separately for each schedule. We then ran multiple seeds using the best learning-rate from the grid search. Mean and standard error of the mean was tabulated for each schedule. The best result for each method, up to a statistical significance level of 0.05 using a paired two-sample t-test, is highlighted in bold. Specific details of the hyper-parameters used for each problem are given in Appendix G. All non-adaptive schedules included a fixed learning-rate warmup, with length following standard practices for the problem. Our step-wise schedule uses a 30-60-90 percent tenthing, following standard practices from ImageNet training. As in the convex experiments, the Refined schedules were calculated by using the linear schedule to generate the initial gradient norm sequence.
Tables 2 & 3 show the results. We break the schedules into two categories, classical and modern. The modern schedules consistently outperform the classical schedules, often by large margins. Although this is common folk-law in deep learning, we are not aware of any existing large-scale experiments establishing this. Our comparison of modern schedules shows a clear hierarchy among the schedules. The Stepwise schedule is dominated by the Cosine schedule, and the Linear schedule matches or outperforms the Cosine schedule on all problems except ViT. The refined schedule further outperforms the Linear schedule on 5 of the 10 problems, but shows mixed results on ViT and RCNN. The refinement process produced degenerate schedules that fail on the CIFAR problems, we discuss this in Appendix 3.5.
Table 2: Classical Schedule Comparison Against the Linear Schedule (lower = better).
Problem Flat 1/t 1/sqrt Offset 1/t Offset 1/sqrt Linear
CIFAR10 $8.04$$\pm.13$ $5.42$$\pm.28$ $6.37$$\pm.41$ $9.23$$\pm.08$ $6.58$$\pm.06$ $\mathbf{4.35}$.$\pm.05$
CIFAR100 $30.43$$\pm.20$ $26.58$$\pm.11$ $29.09$$\pm.16$ $32.80$$\pm.07$ $27.62$$\pm.13$ $\mathbf{22.11}$.$\pm.08$
ImageNet $33.00$$\pm.14$ $26.48$$\pm.06$ $28.35$$\pm.05$ $47.94$$\pm.08$ $27.34$$\pm.06$ $\mathbf{23.11}$.$\pm.07$
IWSLT14 $8.07$$\pm.02$ $7.62$$\pm.01$ $7.52$$\pm.01$ $12.89$$\pm.06$ $8.48$$\pm.01$ $\mathbf{7.10}$.$\pm.01$
GPT $20.20$$\pm.000$ $18.99$$\pm.04$ $19.48$$\pm.02$ $27.85$$\pm.05$ $22.88$$\pm.006$ $\mathbf{18.60}$.$\pm.02$
RoBERTa $4.52$$\pm.005$ $4.25$$\pm.007$ $4.33$$\pm.01$ $5.33$$\pm.02$ $5.15$$\pm.02$ $\mathbf{3.94}$.$\pm.007$
DLRM $\mathbf{20.95}$.$\pm.006$ $47.59$$\pm 6.45$ $45.99$$\pm 5.98$ $\mathbf{20.94}$.$\pm.007$ $20.99$$\pm.009$ $\mathbf{20.94}$.$\pm.006$
MRI $\mathbf{9.00}$.$\pm.04$ $8.91$$\pm.01$ $\mathbf{8.98}$.$\pm.04$ $9.53$$\pm.08$ $9.16$$\pm.05$ $\mathbf{8.88}$.$\pm.02$
ViT $30.11$$\pm.27$ $28.36$$\pm.40$ $28.53$$\pm.15$ $73.84$$\pm 6.08$ $50.36$$\pm 12.39$ $\mathbf{24.82}$.$\pm.31$
RCNN $65.43$$\pm.12$ $63.38$$\pm.05$ $64.13$$\pm.10$ $79.32$$\pm.07$ $69.25$$\pm.07$ $\mathbf{60.98}$.$\pm.02$
Table 3: Modern Schedule Comparison (lower = better).
Problem Stepwise Cosine Linear Refinement
CIFAR10 $4.53$$\pm.03$ $\mathbf{4.27}$.$\pm.04$ $\mathbf{4.35}$.$\pm.05$
CIFAR100 $22.78$$\pm.10$ $22.59$$\pm.09$ $\mathbf{22.11}$.$\pm.08$
ImageNet $23.51$$\pm.07$ $\mathbf{23.10}$.$\pm.06$ $\mathbf{23.11}$.$\pm.07$ $\mathbf{23.12}$$\pm 0.03$
IWSLT14 $7.43$$\pm.01$ $7.17$$\pm.009$ $7.10$$\pm.01$ $\mathbf{6.92}$.$\pm.03$
GPT $19.70$$\pm.03$ $18.65$$\pm.02$ $18.60$$\pm.02$ $\mathbf{18.29}$.$\pm.005$
RoBERTa $4.36$$\pm.01$ $4.07$$\pm.000$ $3.94$$\pm.007$ $\mathbf{3.86}$.$\pm.005$
DLRM $20.95$$\pm.008$ $\mathbf{20.94}$.$\pm.005$ $\mathbf{20.94}$.$\pm.006$ $\mathbf{20.94}$.$\pm.009$
MRI $8.97$$\pm.02$ $8.90$$\pm.03$ $8.88$ ±.$\pm.02$ $\mathbf{8.85}$.$\pm.01$
ViT $26.27$$\pm.33$ $\mathbf{24.56}$.$\pm.15$ $24.82$$\pm.31$ $25.53$$\pm.16$
RCNN $61.76$$\pm.06$ $\mathbf{61.00}$.$\pm.04$ $\mathbf{60.98}$.$\pm.02$ $61.88$$\pm.02$
3.3 Cosine schedule ablations
To further investigate the failure modes of the cosine schedule, we ran a series of shorter duration training runs on CIFAR10. Our premise was that the cosine schedule is heavily over-fit to long training duration computer vision problems. As shown in Table 4, the cosine schedule begins to under-perform the linear decay schedule and the refined schedules when training for less than 30 epochs. In contrast, while the refined schedule also under-performs for longer duration training it has no statistically significant differences from the linear decay schedule for shorter duration training, where they both perform consistently better than the cosine schedule.
Table 4: CIFAR10 Training Time Ablations (Test Error %).
Schedule 1 Epoch 5 Epochs 15 Epochs 30 Epochs
Cosine $35.80$$\pm 0.27$ $15.07$$\pm 0.11$ $7.99$$\pm 0.06$ $\mathbf{6.08}$$\pm 0.03$
Linear Decay $\mathbf{33.84}$$\pm 0.17$ $\mathbf{14.36}$$\pm 0.07$ $\mathbf{7.65}$$\pm 0.07$ $\mathbf{6.12}$$\pm 0.08$
Refined $\mathbf{34.04}$$\pm 0.19$ $\mathbf{14.59}$$\pm 0.08$ $7.88$$\pm 0.04$ $6.24$$\pm 0.06$
Figure 4: Limitations of refinement: if model over-fits, our method produces degenerate schedules.
3.4 Large Language Model size ablations
Figure 5: Normalized learning rate schedules learned on vanilla Transformer-based LLM trained on C4 dataset. Left: schedules obtained when setting $w_{t}\propto\|g_{t}\|^{-2}$. Right: schedules when $w_{t}\propto 1/\|g_{t}\|_{1}$. We find that $\ell_{1}$ norm is much more consistent across model sizes and baseline schedules. Thus, we used $w_{t}\propto 1/\|g_{t}\|_{1}$ weighting from linear baselines to obtained Refined schedule mentioned in Table 5. The x axis is training steps.
Table 5: LLM Ablations (C4 validation loss).
Schedule 117M 284M 1B 3.5B
Cosine $3.089$ $2.891$ $2.729$ $2.631$
Linear Decay $3.087$ $2.888$ $2.725$ $\mathbf{2.625}$
Refined $\mathbf{3.075}$ $\mathbf{2.884}$ $\mathbf{2.722}$ $2.634$
In addition to the results above, we validate our insights on Language Models by performing an additional set of experiments directly comparing to the setting explored in Chinchilla (Hoffmann et al., 2022) where we train a vanilla Transformer model on the C4 dataset (Raffel et al., 2020) for 10k steps with a token batch size of around 524k ($2^{19}$). Starting with Hoffmann et al. (2022), most recent LLMs employ the AdamW optimizer with Linear Warmup and Cosine Decay. We perform a head-to-head comparison of Cosine decay, Linear decay and Refined schedules (the latter is refined from Linear decay). As shown in Table 5, contrary to popular wisdom, Linear decay performs better than Cosine decay across all model sizes. The Refined schedule outperforms Linear Decay for all but the 3.5B sized model.
3.5 Limitations
The cifar10 and cifar100 problems show potential pitfalls from using our approach (Figure 4) when our assumptions break down. On these problems $100\%$ train accuracy is reached near the end, driving the gradient norm sequence to zero. The resulting refined schedule rapidly increases the step size near the end of optimization, which if used results in training divergence. We suggest using a linear decay schedule instead on problems where the gradient norm drops significantly at the end.
4 Discussion
Gradient sequences that slowly decrease or slowly increase after stabilizing are often observed in practice (Figure 2), and the resulting schedules that our framework produces show interesting behavior, resembling polynomial-decay schedules with exponents $p$:
$\eta_{t}\propto\left(1-\frac{t}{T}\right)^{p}.$ (3)
Schedules of this form are already in use, with $p$ typically tuned as a hyper-parameter. From Figure 1, we see that $p<1$ should be used when the gradient norm is decreasing over time, $p>1$ when the gradient norm is increasing, and $p=1$ when the gradient sequence is flat. We use the term polynomial decay for the schedule given in Equation 3, which its name in both PyTorch and Tensorflow, although in the literature schedules of the form $1/t^{\alpha}$ are also sometimes referred to as polynomial schedules. From Figure 2, we can see that many of our test problems have refined schedules that are well-approximated by a polynomial decay sequences, shown in blue.
5 Related work
The Robbins-Monro conditions are the foundation of early learning rate theory (Robbins and Monro, 1951). They advocated for step size sequences where $\sum_{t=1}^{\infty}\eta_{t}=\infty$ and $\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty$. Of the schedules satisfying these conditions, they advocated for schedules with $1/t$ decay as they are asymptotically optimal for twice continuously differentiable and strongly convex functions. This schedule was later shown to be optimal (even non-asymptotically) for strongly convex G-Lipschitz stochastic optimization when appropriate averaging is used (Shamir and Zhang, 2013; Lacoste-Julien et al., 2012; Rakhlin et al., 2012).
Nemirovski et al. (2009) point out that the performance of these step-sizes is extremely poor when the base step size is miss-specified, and advocate $1/\sqrt{t}$ step sizes, which have the advantage of performing asymptotically optimally when strong convexity is not present. The use of $1/\sqrt{t}$ step sizes in combination with averaging was originally developed by Polyak (1990) and Ruppert (1988), and are optimal for the complexity class of G-Lipschitz convex functions with or without stochasticity (Blair, 1985; Nesterov, 2018).
Scheduling for the stochastic quadratic setting has seen significant recent interest. The seminal work of Bach and Moulines (2013) establish that a flat $1/4R$ schedule (for a given problem dependent-constant $R$), independent of the stopping time $T$, gives the optimal $\mathcal{O}(1/T)$ rate for the average iterate. For the last iterate, Ge et al. (2019a) establish that the step-decay schedule’s rate:
$\Theta\left(\frac{d\sigma^{2}}{T}\log T\right),$
Table 6: Polynomial Decay Schedules fit to refined schedules
Problem Warmup Power
DLRM 20% 0.9
GPT 28% 1.0
RoBERTa 45% 0.8
MRI 40% 0.9
RCNN 0% 1.5
ImageNet 1% 3.0
is a multiplicative log factor worse than the mini-max optimal rate. Here $d$ is the dimensionality of the problem and $\sigma$ a bound on the noise. They further show that the any-time last-iterate behavior of SGD is poor, visiting highly sub-optimal function values infinitely often. Pan et al. (2022) study the situation where the full Hessian spectrum is know in advance. In this case, they are able to improve the rate to
$\mathcal{O}\left(\frac{d\sigma^{2}}{T}\log\kappa\right),$
or, under a skewed spectrum assumption, to
$\mathcal{O}\left(\frac{d\sigma^{2}}{T}\right).$
Ge et al. (2019b) show that commonly used schedules such as polynomial decay schedules are sub-optimal in the quadratic setting.
5.1 Non-smooth last iterate convergence
For non-smooth stochastic optimization, Shamir and Zhang (2013) show that the last iterate of SGD on non-smooth, convex problems with a $1/\sqrt{t}$ step size is
$\mathcal{O}\left(\frac{\log(n)}{\sqrt{n}}\right).$
Harvey et al. (2019) show a corresponding lower bound, showing that the log factor is unavoidable when using $1/\sqrt{t}$ step-sizes.
Jain et al. (2019) were the first to show that the log-factor can be removed if the horizon is known in advance. Their approach uses a sequence of step sizes:
$\gamma_{t}=\frac{D2^{-i}}{G\sqrt{T+1}},$
where $i$ is an epoch counter, for a sequence of epochs each half the length of the preceding epoch. This step size sequence gives a convergence rate for the last iterate of:
$E\left[f(x_{T})-f_{*}\right]\leq\frac{15GD}{\sqrt{T}}.$
For more general conversions from regret bounds to last iterate guarantees, Cutkosky (2019) provides an alternative conversion based upon evaluating gradients at running averages of iterates and achieves the same last-iterate convergence rate as Theorem 1. However, our method is directly interpretable in terms of learning rate schedules to yield immediate practical guidance while the technique of Cutkosky (2019) is a more opaque transformation of the base algorithm. Nevertheless, there are interesting contrasts: the result of Cutkosky (2019) is ”anytime” in the sense that each individual iterate comes with an optimal convergence guarantee. In our case, only the last iterate has an optimal guarantee (in fact, Zamani and Glineur (2023) show that this restriction is necessary for any scheme based on learning rate schedules). Understanding the differences in these approaches is a valuable future direction.
5.2 Schedules in Applications
Learning rate schedules used in applications usually fall within a few standard classes supported by major frameworks. To survey the use of schedulers in open source software, we performed a GitHub search on each of the schedulers in the PyTorch Library. Table 7 shows that the polynomial decay schedule that we advocate is currently the least popular scheduler in PyTorch.
PyTorch Scheduler Github Files (in thousands)
ReduceLROnPlateau 105
StepLR 101
MultiStepLR 37.9
CosineAnnealingLR 37.1
ExponentialLR 16
OneCycleLR 14.9
CosineAnnealingWarmRestarts 10.9
CyclicLR 9.1
LinearLR 5.9
ConstantLR 3.6
MultiplicativeLR 2.6
PolynomialLR 1.3
Table 7: Popularity of standard learning rate schedules
The reduce-on-plateau, and its fixed length sister, the step-wise schedule, popularized by Krizhevsky et al. (2012) for use in image classification, is considered a benchmark method. They justify it heuristically (“The heuristic which we followed was to divide the learning rate by 10 when the validation error rate stopped improving with the current learning rate.”). Due to its use in highly influential papers as well as its use in tutorials, it is by far the most commonly used schedule (Table 7).
Many early works use step-sizes satisfying the Robbins-Monro conditions, such as step sizes proportional to $1/(c+\alpha t)$, for some $c$ and $\alpha$, as advocated in the influential work of Bottou (1991, 2012). These step sizes produce large values for small $t$ and so are often clipped (using $\min(\beta,\cdot)$ for some $\beta$).
Schedules involving restarting (also called cycling) are also widely used, particularly on computer vision problems where overfitting can be an issue (Smith and Topin, 2018). Loshchilov and Hutter (2017b) use restarting together with the cosine annealing schedule:
$\eta_{t}=\eta_{\min}+\left(\eta_{\max}-\eta_{\min}\right)\left(1+\cos\left(\frac{t}{T}\pi\right)\right),$ (4)
This schedule is remarkably effective in practice for image classification problems even without restarts; as far as the authors are aware, for ResNet-50 ImageNet training, no schedule has been demonstrated that out-performs it. It is less commonly used in other problem domains. The motivation for the use of cosine function is unstated, and there is no theory to support it. The reasons for the success of the schedule are not currently understood.
The only large-scale comparison of schedulers for deep learning optimization that we are aware of is the work of Schmidt et al. (2021). They compare on a smaller set of problems than our comparison, although across a range of optimizers, whereas we focus on just Adam & SGD.
5.3 Any-time Adaptive Schedules
There has been significant work on any-time learning rate adaptation that makes use of gradient or loss information to adapt the learning rate. Schedules in the AdaGradNorm class are backed by strong theory (Wu et al., 2020; Streeter and McMahan, 2010; Duchi et al., 2011; Ward et al., 2019; Li and Orabona, 2019). They use step sizes of the form:
$\eta_{t}\propto\frac{1}{\sqrt{\sum_{t=1}^{T}\left\|g_{t}\right\|^{2}}}$
They have not seen significant adoption in practice, likely due to the intrinsic disadvantages of any-time $O(1/\sqrt{t})$-style schedules as already discussed. It is also not clear how to apply this schedule on top of existing high-performance optimizers such as Adam.
Methods based on hyper-gradients (i.e. gradients of the learning rate) are another heavily studied class of any-time adaptive methods. They use recent gradient evaluations and other local information to update the learning rate at each step of training (Almeida et al., 1999; Franceschi et al., 2017; Bengio, 2000; Domke, 2012; Pedregosa, 2016; Baydin et al., 2018; Feurer and Hutter, 2019; Donini et al., 2020; Chandra et al., 2022). The hyper-gradient can also be used in a more cautious manor to define conditions under which the learning rate should be decreased (Pflug, 1983, 1988; Lang et al., 2019; Zhang et al., 2020).
Methods based on loss minimization aim to choose a learning rate that greedily minimize the loss, either at each step, or based on aggregate statistics that are either produced in stages or by smoothing. They include methods that target the train loss or loss on held out validation data (Xu et al., 2019; Jin et al., 2021; Kim et al., 2021).
Methods that query additional points, such as line search methods, are not scheduling methods per se but can be used in place of scheduling (Mahsereci and Hennig, 2017; Iyer et al., 2021). Due to the additional implementation complexity and the cost of their use, they also have not seen widespread adoption for deep learning.
Conclusion
Our experimental results strongly support the use of the linear decay schedule as a default baseline for deep learning optimization. Applying refinement on top of an initial linear decay run gives a schedule showing improvements on 5 out of 10 deep learning problems and 5 of 8 convex problems. We develop extensive theory supporting the linear decay schedule and our novel refinement technique, establishing that their SOTA performance is well-supported both in theory and practice.
Appendix A All-Tail Summation Bound
In this section we develop some key Lemmas that underlie our results. The starting point of this development is the following result from Orabona (2020); Lin et al. (2016). We present the result and proof below for completeness, and then provide an improved version of the bound (Lemma 5) that is used to derive our analytical results.
A.1 Existing bound
Lemma 4
Let $q_{t}\geq 0$, and let $\eta_{t}$ be a positive non-increasing sequence. Then:
$\eta_{T}q_{T}\leq\frac{1}{T}\sum_{t=1}^{T}\eta_{t}q_{t}+\sum_{k=1}^{T-1}\frac{1}{k(k+1)}\sum_{t=k+1}^{T}\eta_{t}\left(q_{t}-q_{k}\right),$
Proof  We present below a version of the proof that’s written to move the single inequality application in the whole proof to the last step. Define $S_{k}=\frac{1}{k}\sum_{t=T-k+1}^{T}\eta_{t}q_{t}$.
Then
$kS_{k}$ $=(k+1)S_{k+1}-\eta_{T-k}q_{T-k}$
$=kS_{k+1}+S_{k+1}-\eta_{T-k}q_{T-k}$
$=kS_{k+1}+\frac{1}{k+1}\sum_{t=T-k}^{T}\left(\eta_{t}q_{t}-\eta_{T-k}q_{T-k}\right).$
Dividing through by $k$:
$S_{k}=S_{k+1}+\frac{1}{k(k+1)}\sum_{t=T-k}^{T}\left(\eta_{t}q_{t}-\eta_{T-k}q_{T-k}\right).$
Unrolling, we have:
$S_{1}=S_{T}+\sum_{k=1}^{T-1}\frac{1}{k(k+1)}\sum_{t=T-k}^{T}\left(\eta_{t}q_{t}-\eta_{T-k}q_{T-k}\right).$
Now we use $S_{1}=\eta_{T}q_{T}$. Note that the first entry in that sum is zero so we may shift the indexing to start at $t=T-k+1$. Giving:
$\eta_{T}q_{T}$ $=S_{T}+\sum_{k=1}^{T-1}\frac{1}{k(k+1)}\sum_{t=T-k+1}^{T}\left(\eta_{t}q_{t}-\eta_{T-k}q_{T-k}\right)$
$\leq S_{T}+\sum_{k=1}^{T-1}\frac{1}{k(k+1)}\sum_{t=T-k+1}^{T}\eta_{t}\left(q_{t}-q_{T-k}\right).$
Where the final step uses the fact that $\eta_{t}$ is decreasing. Plugging in the definition of $S_{T}$, and substituting $k=T-k$ to simplify gives the result.
A.2 Improved expression
Lemma 5
Let $q_{t}$ be any sequence, and let $w_{t}$ be a positive sequence. Then:
$q_{T}$ $=\frac{1}{\sum_{t=1}^{T}w_{t}}\sum_{t=1}^{T}w_{t}q_{t}+\sum_{k=1}^{T-1}\frac{w_{k}}{\sum_{t=k+1}^{T}w_{t}}\left(\frac{1}{\sum_{t=k}^{T}w_{t}}\sum_{t=k}^{T}w_{t}\left(q_{t}-q_{k}\right)\right)$
$=\frac{1}{w_{1:T}}\sum_{t=1}^{T}w_{t}q_{t}+\sum_{k=1}^{T-1}\left(\frac{1}{w_{k+1:T}}-\frac{1}{w_{k:T}}\right)\sum_{t=k}^{T}w_{t}\left(q_{t}-q_{k}\right).$
Proof  Define
$S_{k}=\frac{1}{\sum_{t=T-k+1}^{T}w_{t}}\sum_{t=T-k+1}^{T}w_{t}q_{t}.$
Note that with this definition:
$S_{1}=\frac{1}{\sum_{t=T}^{T}w_{t}}\sum_{t=T}^{T}w_{t}q_{t}=q_{T},$
and $S_{T}$ is the full sum:
$S_{T}=\frac{1}{\sum_{t=1}^{T}w_{t}}\sum_{t=1}^{T}w_{t}q_{t}.$
The difference from the weighting used in the the Lemma above: we normalized by the sum of the step sizes rather than $k$. We get the following expansion:
$\left(\sum_{t=T-k+1}^{T}w_{t}\right)S_{k}$ $=\sum_{t=T-k+1}^{T}w_{t}q_{t}$
$=\sum_{t=T-k}^{T}w_{t}q_{t}-w_{T-k}q_{T-k}$
$=\left(\sum_{t=T-k}^{T}w_{t}\right)S_{k+1}-w_{T-k}q_{T-k}$
$=\left(\sum_{t=T-k+1}^{T}w_{t}\right)S_{k+1}+\left(w_{T-k}S_{k+1}-w_{T-k}q_{T-k}\right).$
So dividing through by $\sum_{t=T-k+1}^{T}w_{t}$:
$S_{k}=S_{k+1}+\frac{w_{T-k}}{\sum_{t=T-k+1}^{T}w_{t}}\left(S_{k+1}-q_{T-k}\right).$
Unrolling
$S_{1}=S_{T}+\sum_{k=1}^{T-1}\frac{w_{T-k}}{\sum_{t=T-k+1}^{T}w_{t}}\left(S_{k+1}-q_{T-k}\right).$
Note that, plugging in $S_{k+1}$:
$S_{k+1}-q_{T-k}$ $=\frac{1}{\sum_{t=T-k}^{T}w_{t}}\sum_{t=T-k}^{T}w_{t}q_{t}-q_{T-k}$
$=\frac{1}{\sum_{t=T-k}^{T}w_{t}}\sum_{t=T-k}^{T}w_{t}\left(q_{t}-q_{T-k}\right).$
So we have:
$q_{T}=\frac{1}{\sum_{t=1}^{T}w_{t}}\sum_{t=1}^{T}w_{t}q_{t}+\sum_{k=1}^{T-1}\frac{w_{T-k}}{\sum_{t=T-k+1}^{T}w_{t}}\left(\frac{1}{\sum_{t=T-k}^{T}w_{t}}\sum_{t=T-k}^{T}w_{t}\left(q_{t}-q_{T-k}\right)\right).$
Finally, we make the simple change of variables $k\to T-k$ to yield the result.
Remark 6
We will typically use this result by setting $q_{t}=f(x_{t})-f_{*}$ in order to bound $q_{T}=f(x_{T})-f_{*}$. By using a weighting sequence $w_{t}$ elsewhere, we are able to remove the $w_{T}$ weight from in front of the $q_{T}$ term (in contrast to Lemma 4). This is crucial, as we want to be able to analyze the situation in which $w_{t}$ drops extremely small near the end, and yet still bound $q_{T}=f(x_{T})-f_{*}$, but if $q_{T}$ is weighted by $w_{T}$ we will get very loose bounds when $w_{T}$ is small. Notice also that we have an equality instead of an inequality, and we have not had to impose the requirement that $w$ be a non-increasing sequence.
Appendix B Proof of Theorem  1
Before proving Theorem 1, we need the following important Lemma:
Lemma 7
Suppose $z_{1},\dots,z_{T}$ is an arbitrary sequence of vectors. Let $w_{1},\dots,w_{T}$ be an arbitrary sequence of positive numbers. Define the sequence $x_{1},\dots,x_{T}$ recursively by $x_{1}=z_{1}$ and:
$x_{t}=w_{t:T}\left(\frac{z_{t}}{w_{1:T}}+\sum_{p=1}^{t-1}x_{p}\left(\frac{1}{w_{p+1:T}}-\frac{1}{w_{p:T}}\right)\right).$
Suppose $g_{t}$ are random variables. with $\mathbb{E}[g_{t}|x_{1},\dots,x_{t}]\in\partial f(x_{t})$ for some convex $f$. Then:
$\mathbb{E}[f(x_{T})-f(u)]$ $\leq\mathbb{E}\left[\frac{1}{w_{1:T}}\sum_{t=1}^{T}w_{t}\langle g_{t},z_{t}-u\rangle\right].$
Proof  Let us set $q_{t}=f(x_{t})-f(u)$. Then we have $\mathbb{E}[q_{t}]\leq\mathbb{E}[\langle g_{t},x_{t}-u\rangle]$ and $\mathbb{E}[q_{t}-q_{k}]\leq\mathbb{E}[\langle g_{t},x_{t}-x_{k}\rangle]$. Then, Lemma 5 implies:
$\mathbb{E}[q_{T}]$ $\leq\mathbb{E}\left[\frac{1}{w_{1:T}}\sum_{t=1}^{T}w_{t}\langle g_{t},x_{t}-u\rangle+\sum_{k=1}^{T-1}\left(\frac{1}{w_{k+1:T}}-\frac{1}{w_{k:T}}\right)\sum_{t=k}^{T}w_{t}\langle g_{t},x_{t}-x_{k}\rangle\right].$
Now, let us find the coefficient of $\langle g_{t},x_{t}\rangle$ in the above expression. This is:
$\frac{w_{t}}{w_{1:T}}+\left[\sum_{k=1}^{t}\left(\frac{1}{w_{k+1:T}}-\frac{1}{w_{k:T}}\right)w_{t}\right]-w_{t}\left(\frac{1}{w_{t+1:T}}-\frac{1}{w_{t:T}}\right)$
$=\frac{w_{t}}{w_{1:T}}+\sum_{k=1}^{t-1}\left(\frac{1}{w_{k+1:T}}-\frac{1}{w_{k:T}}\right)w_{t}$
$=\frac{w_{t}}{w_{t:T}}.$
Next, for $p<t$, the coefficient of $\langle g_{t},x_{p}\rangle$ is:
$-\left(\frac{1}{w_{p+1:T}}-\frac{1}{w_{p:T}}\right)w_{t}.$
And for $p>t$, the coefficient of $\langle g_{t},x_{p}\rangle$ is zero. Finally the coefficient of $\langle g_{t},u\rangle$ is $-\frac{w_{t}}{w_{1:T}}$.
Putting this all together, we can rearrange the expression as follows:
$\mathbb{E}[q_{T}]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}\left\langle g_{t},\frac{w_{t}}{w_{t:T}}x_{t}-\left(\frac{w_{t}u}{w_{1:T}}+\sum_{p=1}^{t-1}w_{t}x_{t}\left(\frac{1}{w_{p+1:T}}-\frac{1}{w_{p:T}}\right)\right)\right\rangle\right]$
$=\mathbb{E}\left[\sum_{t=1}^{T}\left\langle\frac{w_{t}}{w_{t:T}}g_{t},x_{t}-w_{t:T}\left(\frac{u}{w_{1:T}}+\sum_{p=1}^{t-1}x_{p}\left(\frac{1}{w_{p+1:T}}-\frac{1}{w_{p:T}}\right)\right)\right\rangle\right].$
Now, given an arbitrary sequence $z_{1},\dots,z_{T}$, define $x_{t}$ recursively by:
$x_{t}=w_{t:T}\left(\frac{z_{t}}{w_{1:T}}+\sum_{p=1}^{t-1}x_{p}\left(\frac{1}{w_{p+1:T}}-\frac{1}{w_{p:T}}\right)\right).$
Then we have:
$\mathbb{E}[q_{T}]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}\left\langle\frac{w_{t}}{w_{t:T}}g_{t},x_{t}-w_{t:T}\left(\frac{u}{w_{1:T}}+\sum_{p=1}^{t-1}x_{p}\left(\frac{1}{w_{p+1:T}}-\frac{1}{w_{p:T}}\right)\right)\right\rangle\right]$
$=\mathbb{E}\left[\sum_{t=1}^{T}\frac{w_{t}}{w_{1:T}}\langle g_{t},z_{t}-u\rangle\right].$
Now, we are finally ready to prove Theorem 1:
See 1
Proof  Let’s define $\hat{x}_{1}=z_{t}$ and recursively set:
$\hat{x}_{t}=w_{t:T}\left(\frac{z_{t}}{w_{1:T}}+\sum_{p=1}^{t-1}\hat{x}_{p}\left(\frac{1}{w_{p+1:T}}-\frac{1}{w_{p:T}}\right)\right).$
Then, Lemma 7 shows that $\mathbb{E}[f(\hat{x}_{T})-f(u)]\leq\mathbb{E}\left[\frac{1}{w_{1:T}}\sum_{t=1}^{T}w_{t}\langle g_{t},z_{t}-u\rangle\right]$. So, it suffices to show that $x_{t}=\hat{x}_{t}$ for all $t$. In turn, since $\hat{x}_{1}=z_{1}=x_{1}$, it suffices to show $\hat{x}_{t+1}-\hat{x}_{t}=\frac{w_{t+1:T}}{w_{1:T}}\Delta_{t}=x_{t+1}-x_{t}$ for all $t$.
To this end, let’s do some calculation. First:
$\hat{x}_{t}$ $=\frac{w_{t}}{w_{t:T}}\hat{x}_{t}+\frac{w_{t+1:T}}{w_{t:T}}\hat{x}_{t}$
$=\frac{w_{t}}{w_{t:T}}\hat{x}_{t}+w_{t+t:T}\left(\frac{z_{t}}{w_{1:T}}+\sum_{p=1}^{t-1}\hat{x}_{p}\left(\frac{1}{w_{p+1:T}}-\frac{1}{w_{p:T}}\right)\right).$
With this expression, we have:
$\hat{x}_{t+1}-\hat{x}_{t}$ $=\hat{x}_{t+1}-w_{t+1:T}\left(\frac{z_{t+1}}{w_{1:T}}+\sum_{p=1}^{t}x_{p}\left(\frac{1}{w_{p+1:T}}-\frac{1}{w_{p:T}}\right)\right)-\frac{w_{t}}{w_{t:T}}\hat{x}_{t}$
$=w_{t+1:T}\left(\frac{z_{t+1}}{w_{1:T}}+\sum_{p=1}^{t}\hat{x}_{p}\left(\frac{1}{w_{p+1:T}}-\frac{1}{w_{p:T}}\right)\right)$
$\qquad-w_{t+1:T}\left(\frac{z_{t}}{w_{1:T}}+\sum_{p=1}^{t-1}x_{p}\left(\frac{1}{w_{p+1:T}}-\frac{1}{w_{p:T}}\right)\right)-\frac{w_{t}}{w_{t:T}}\hat{x}_{t}$
$=\frac{w_{t+1:T}(z_{t+1}-z_{t})}{w_{1:T}}+w_{t+1:T}\hat{x}_{t}\left(\frac{1}{w_{t+1:T}}-\frac{1}{w_{t:T}}\right)-\frac{w_{t}}{w_{t:T}}\hat{x}_{t}$
$=\frac{w_{t+1:T}}{w_{1:T}}\Delta_{t}+\hat{x}_{t}\left(1-\frac{w_{t+1:T}+w_{t}}{w_{t:T}}\right)$
$=\frac{w_{t+1:T}}{w_{1:T}}\Delta_{t}.$
Appendix C Proof of Theorem 3
See 3
Proof  First, observe that with $\Delta_{t}=-w_{t}\eta_{t}$, we have $x_{t+1}=x_{t}+\frac{w_{t+1:T}}{w_{1:T}}\Delta_{t}$. Therefore with $z_{1}=x_{1}$ and $z_{t+1}=z_{t}+\Delta_{t}$, Theorem 1 implies:
$\mathbb{E}[f(x_{T})-f(u)]$ $\leq\mathbb{E}\left[\frac{1}{w_{1:T}}\sum_{t=1}^{T}\langle w_{t}g_{t},z_{t}-u\rangle\right].$
Next, observe that $z_{t}$ is simply online gradient descent with learning rate 1 acting on the loss vectors $w_{t}z_{t}$. Standard analysis (Zinkevich, 2003) shows:
$\sum_{t=1}^{T}\langle w_{t}g_{t},z_{t}-u\rangle$ $=\frac{\|z_{1}-u\|^{2}}{2}-\frac{\|z_{T+1}-u\|^{2}}{2}+\sum_{t=1}^{T}\frac{w_{t}^{2}\|g_{t}\|^{2}}{2}.$
This immediately implies the first part of the Theorem. Next, we need to solve for the minimizing values of $w_{t}$. To do this, we take the logarithm of the expression $\frac{1}{2w_{1:T}}\left(\|x_{1}-u\|^{2}+\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2}\right)$ and differentiate:
$\frac{\partial}{\partial w_{k}}\log\left[\frac{1}{2w_{1:T}}\left(\|x_{1}-u\|^{2}+\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2}\right)\right]$ $=\frac{2w_{k}\|g_{k}\|^{2}}{\|x_{1}-u\|^{2}+\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2}}-\frac{1}{w_{1:T}}.$
We set this equal to zero to solve for the optimal $w_{k}$:
$w_{k}$ $=\|g_{k}\|^{-2}\frac{\|x_{1}-u\|^{2}+\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2}}{2w_{1:T}}\triangleq\lambda\|g_{k}\|^{-2},$
where we have defined $\lambda=\frac{\|x_{1}-u\|^{2}+\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2}}{2w_{1:T}}$, which does not depend on $k$. That is, the optimal $w_{k}$ value is proportional to $\|g_{k}\|^{-2}$. With this expression, we have:
$\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2}$ $=\lambda^{2}\sum_{t=1}^{T}\|g_{t}\|^{-2}$
$w_{1:T}$ $=\lambda\sum_{t=1}^{T}\|g_{t}\|^{-2}.$
So, let us now solve for $\lambda$ by plugging in these values:
$\lambda$ $=\frac{\|x_{1}-u\|^{2}+\sum_{t=1}^{T}w_{t}^{2}\|g_{t}\|^{2}}{2w_{1:T}}$
$\lambda$ $=\frac{\|x_{1}-u\|^{2}+\lambda^{2}\sum_{t=1}^{T}\|g_{t}\|^{-2}}{2\lambda\sum_{t=1}^{T}\|g_{t}\|^{-2}}$
$\lambda$ $=\frac{\|x_{1}-u\|^{2}}{2\lambda\sum_{t=1}^{T}\|g_{t}\|^{-2}}+\frac{\lambda}{2}$
$\lambda$ $=\frac{\|x_{1}-u\|}{\sqrt{\sum_{t=1}^{T}\|g_{t}\|^{-2}}}.$
This in turn implies the claimed optimal value for $w_{k}$.
Appendix D Proof of Theorem 9
Proof  First, observe that for any solution to the desired identity $\eta_{t}=\frac{w_{t}w_{t:1:T}}{w_{1:T}}$, replacing $w$ with $c\cdot w$ for some constant $c$ will yield a solution for $\eta$ replaced with $c\cdot\eta$. Thus, it suffices to consider the possibility that $\max_{t}\eta_{t}=1$.
The intuition for the rest of the proof is the following: Given the value for $w_{1}$, we can solve for $w_{2:T}$ using the equation $\frac{w_{t}w_{2:T}}{w_{1}+w_{2:T}}=\eta_{1}$. This in turn provides us with the value of $w_{1:T}$. Now, given $w_{k:T}$ for any $k$ we can solve for $w_{k}$ using the equation $\eta_{k}=\frac{w_{k}w_{k+1:T}}{w_{1:T}}=\frac{w_{k}(w_{k:T}-w_{k})}{w_{1:T}}$. Thus we may recursively compute all the values of $w$. Each of these steps requires solving a quadratic equation. We simply choose an initial $w_{1}$ so as to ensure that all the following quadratic equations have non-negative real roots.
Specifically, set $w_{1}=\frac{2^{2T}+\sqrt{2^{4T}-4\eta_{1}}}{2}$ and define $s_{1}=2^{2T}$. Then, recursively define for $t=2,\dots,T-1$:
$s_{t}$ $=s_{t-1}-w_{t-1}$
$w_{t}$ $=\frac{s_{t}-\sqrt{s_{t}^{2}-4s_{1}\eta_{t}}}{2}$
and set $w_{T}=s_{T}=s_{T-1}-w_{T-1}$. Notice that these choices satisfy:
$w_{t}^{2}-s_{t}w_{t}+s_{1}\eta_{t}=0$
so that if we could establish (1) that all $w_{t}\geq 0$ and (2) $s_{t}=w_{t:T}$, then we would have:
$\frac{w_{t}w_{t+1:T}}{w_{1:T}}$ $=\frac{w_{t}(w_{t:T}-w_{t})}{w_{1:T}}$
$=\frac{w_{t}(s_{t}-w_{t})}{s_{1}}$
$=\frac{s_{t}w_{t}-w_{t}^{2}}{s_{1}}$
$=\eta_{t}$
as desired.
Let us first prove (1): all $w_{t}\geq 0$. For $t\leq T-1$, this will hold if $s_{t}^{2}-4s_{1}\eta_{t}>0$ To establish this, we will first show that $s_{t}\geq\frac{s_{1}}{2^{t-1}}$ for all $t\leq T-1$. If this holds, then we have:
$s_{t}^{2}\geq\frac{s_{1}^{2}}{2^{2t-2}}\geq\frac{s_{1}2^{2T}}{2^{2t-2}}\geq 4s_{1}\geq 4s_{1}\eta_{t}$
for $t\leq T-1$, where we have used our assumption $\eta_{t}\leq 1$.
So, we now establish $s_{t}\geq\frac{s_{1}}{2^{t-1}}$ by induction for $t\leq T-1$. The statement is clear for $t=1$. Suppose it holds for for all $t\leq k$ for some $k\leq T-2$. Then we have:
$s_{k+1}$ $=s_{k}-w_{k}$
$=\frac{s_{k}+\sqrt{s_{k}^{2}-4s_{1}\eta_{k}}}{2}$
$\geq\frac{s_{k}}{2}$
$\geq\frac{s_{1}}{2^{k-1+1}},$
which establishes the claim. Therefore for $t\leq T-1$, $w_{t}$ are non-negative real numbers. Finally, we have:
$w_{T-1}$ $=\frac{s_{T-1}+\sqrt{s_{T-1}^{2}-4s_{1}\eta_{t}}}{2}\leq s_{T-1},$
so that $w_{T}=s_{T}=s_{T-1}-w_{T-1}\geq 0$. Thus $w_{t}\geq 0$ for all $t$.
Next, to show (2): $s_{t}=w_{t:T}$. This is nearly immediate. By definition we have $w_{T}=s_{T}$. Suppose $s_{k}=w_{t:T}$ for some $t\geq 2$. Then $s_{t}=s_{t-1}-w_{t-1}$ so that $s_{t-1}=s_{t}+w_{t-1}=w_{t-1:T}$. Thus by induction we have $s_{t}=w_{t:T}$ for all $t$, which is the last thing we needed to show.
Appendix E Schedules for Per-Coordinate Updates
Many popular optimization algorithms in use today like Adam (Kingma and Ba, 2015) and its variants employ per-coordinate updates: the update $\Delta_{t}$ is not proportional to the gradient $g_{t}$ but instead scales each coordinate of $g_{t}$ by an adaptively chosen value. In this section we propose an approximately optimal schedule for such methods.
Theorem 8
Suppose $\Delta_{t}=z_{t+1}-z_{t}=-w_{t}\cdot(\eta_{t}\odot g_{t})$ where $\eta_{t}\in\mathbb{R}^{d}$ is a vector of learning rates and $\odot$ indicates coordinate-wise product. Set $x_{t+1}=x_{t}-\frac{w_{t+1:T}}{w_{1:T}}\Delta_{t}$. Define the quantity $R$ by:
$R=\sqrt{\sum_{t=1}^{T}\sum_{i=1}^{d}\frac{(z_{t,i}-u_{i})^{2}-(z_{t+1,i}-u_{i})^{2}}{\eta_{t,i}}}$
Then we have:
$\mathbb{E}[f(x_{T})-f(u)]$ $\leq\mathbb{E}\left[\frac{1}{w_{1:T}}\left(\frac{R^{2}}{2}+\sum_{t=1}^{T}w_{t}^{2}\sum_{i=1}^{d}\eta_{t,i}g_{t,i}^{2}\right)\right].$
Moreover, for any given fixed values for $R$ and $g_{t,i}^{2}$, the expression $\frac{1}{w_{1:T}}\left(\frac{R^{2}}{2}+\sum_{t=1}^{T}w_{t}^{2}\sum_{i=1}^{d}\eta_{t,i}^{2}g_{t,i}^{2}\right)$ is minimized by setting $w_{t}$ as below:
$w_{t}$ $=\frac{R}{\sqrt{\sum_{t=1}^{T}\left(\sum_{i=1}^{d}\eta_{t,i}g_{t,i}^{2}\right)^{-1}}}\cdot\left(\sum_{i=1}^{d}\eta_{t,i}g_{t,i}^{2}\right)^{-1}.$
As an example of how to use this result, let us suppose we are employing the Adam optimizer, and let us also also ignore the presence of momentum when computing the weights (so really, these will be the weights for the RMSProp optimizer). In this case, $\eta_{t,i}\propto\frac{1}{\sqrt{v_{t,i}}}$ where $v_{t,i}$ is the exponential average of the squared gradients. Thus, we get:
$w_{t}=\lambda\left(\sum_{i=1}^{d}\frac{g_{t,i}^{2}}{\sqrt{v_{t,i}}}\right)^{-1}$
for some $\lambda$. Then, we use a learning rate schedule of $\frac{w_{t}w_{t+1:T}}{w_{1:T}}$.
That is, the corresponding procedure to Algorithm 1 is given by Algorithm 2:
1:Input: $G=\left(G_{t}=\sum_{i=1}^{d}\frac{g_{t,i}^{2}}{\sqrt{v_{t,i}}}\right)$ length $T$ sequence of weighted gradient norms from Adam optimizer, smoothing hyper-parameter $\tau>0$
[…]
3:Define $w_{t}=\hat{G}_{t}^{-1}$
4:For each $t$, let:
5:
$\eta_{t}=w_{t}\sum_{p=t+1}^{T}w_{p}$
6:Return normalized schedule $\eta/\max(\eta)$
Algorithm 2 Schedule Refinement for Adam
In practice, however, recording the value $w_{t}\propto\left(\sum_{i=}^{d}\frac{g_{t,i}^{2}}{\sqrt{v_{t,i}}}\right)^{-1}$ may be difficult (for example, in Pytorch, the Adam implementation is primarily C code rather than Python, which makes it substantially more involved to modify). However, by inspecting the formula for $w_{t}$, we can see that it is likely to be an interpolation between the $1/\|g_{t}\|_{2}^{2}$ and $1/\|g_{t}\|_{1}$ (the L1 norm is not squared). The intuition behind this is that if $v_{t,i}$ has a minimum value of $|g_{t,i}|$. With this minimum value, $w_{t}\propto 1/\|g_{t}\|_{1}$. On the other hand if all $v_{t,i}$ are the same constant, then $w_{t}\propto 1/\|g_{t}\|_{2}^{2}$. In practice we expect behavior closer to the first case and recommend using $w_{t}\propto 1/\|g_{t}\|_{1}$.
Proof [proof of Theorem 8] By standard online gradient descent analysis, we have:
$\langle w_{t}\cdot g_{t},z_{t}-u\rangle$ $=\sum_{i=1}^{d}w_{t}g_{t,i}(z_{t,i}-u_{i})$
$=\sum_{i=1}^{d}\frac{(z_{t,i}-u_{i})^{2}}{2\eta_{t,i}}-\frac{(z_{t+1,i}-u_{i})^{2}}{2\eta_{t,i}}+\frac{\eta_{t,i}}{2}w_{t}^{2}g_{t,i}^{2}.$
Summing this over $t$ from $1$ to $T$, we get
$\sum_{t=1}^{T}\langle w_{t}\cdot g_{t},z_{t}-u\rangle$ $=\sum_{t=1}^{T}\sum_{i=1}^{d}\frac{(z_{t,i}-u_{i})^{2}}{2\eta_{t,i}}-\frac{(z_{t+1,i}-u_{i})^{2}}{2\eta_{t,i}}+\sum_{t=1}^{T}w_{t}^{2}\sum_{i=1}^{d}\frac{\eta_{t,i}g_{t,i}^{2}}{2}$
$=\frac{R^{2}}{2}+\sum_{t=1}^{T}\frac{w_{t}^{2}}{2}\sum_{i=1}^{d}\eta_{t,i}g_{t,i}^{2}.$
The first part of the result now follows from Theorem 1.
Next, we again take the logarithm, differentiate and set equal to zero:
00 0 $=\frac{\partial}{\partial w_{k}}\log\left(\frac{1}{w_{1:T}}\left(\frac{R^{2}}{2}+\sum_{t=1}^{T}\frac{w_{t}^{2}}{2}\sum_{i=1}^{d}\eta_{t,i}g_{t,i}^{2}\right)\right)$
$=\frac{2w_{k}\sum_{i=1}^{d}\eta_{k,i}^{2}g_{k,i}^{2}}{R^{2}+\sum_{t=1}^{T}w_{t}^{2}\sum_{i=1}^{d}\eta_{t,i}g_{t,i}^{2}}-\frac{1}{w_{1:T}}.$
Rearranging, we obtain
$w_{k}$ $=\frac{\left(\sum_{i=1}^{d}\eta_{k,i}^{2}g_{k,i}^{2}\right)^{-1}\left(R^{2}+\sum_{t=1}^{T}w_{t}^{2}\sum_{i=1}^{d}\eta_{t,i}g_{t,i}^{2}\right)}{2w_{1:T}}$
$\triangleq\lambda\left(\sum_{i=1}^{d}\eta_{k,i}g_{k,i}^{2}\right)^{-1},$
where we have collected the non $k$-dependent terms into $\lambda=\frac{R^{2}+\sum_{t=1}^{T}w_{t}^{2}\sum_{i=1}^{d}\eta_{t,i}g_{t,i}^{2}}{2w_{1:T}}$. Now we solve for $\lambda$:
$\lambda$ $=\frac{R^{2}+\sum_{t=1}^{T}w_{t}^{2}\sum_{i=1}^{d}\eta_{t,i}g_{t,i}^{2}}{2w_{1:T}}$
$=\frac{R^{2}+\lambda^{2}\sum_{t=1}^{T}\left(\sum_{i=1}^{d}\eta_{t,i}g_{t,i}^{2}\right)^{-1}}{2\lambda\sum_{t=1}^{T}\left(\sum_{i=1}^{d}\eta_{t,i}g_{t,i}^{2}\right)^{-1}}$
$=\frac{R^{2}}{2\lambda\sum_{t=1}^{T}\left(\sum_{i=1}^{d}\eta_{t,i}g_{t,i}^{2}\right)^{-1}}+\frac{\lambda}{2}$
$=\frac{R}{\sqrt{\sum_{t=1}^{T}\left(\sum_{i=1}^{d}\eta_{t,i}g_{t,i}^{2}\right)^{-1}}}.$
Putting $w_{t}=\lambda\left(\sum_{i=1}^{d}\eta_{t,i}g_{t,i}^{2}\right)^{-1}$ completes the proof.
Appendix F Explicit bounds for arbitrary schedules
Section 2 develops a framework that suggests using learning rates of the form $\eta_{t}=\frac{w_{t}w_{t+1:T}}{w_{1:T}}$ and in Theorem 3 we provided a simple closed-form solution for the optimal values of $w_{t}$. Given this apparently restricted form of $\eta_{t}$, it is natural to ask if this restriction is real: that is, can every schedule $\eta_{1},\dots,\eta_{T-1}$ be represented using some weights $w_{t}$? Theorem 9 shows that the answer is “yes”:
Theorem 9
Let $\eta_{1},\dots,\eta_{T-1}$ be a sequence of non-negative numbers with $\eta_{1}\geq 0$. Then there is a sequence of non-negative weights $w_{1},\dots,w_{T}$ such that $\eta_{t}=\frac{w_{t}w_{t+1:T}}{w_{1:T}}$ for all $t$.
This Theorem shows that by optimizing the weights $w_{t}$, we are in some sense also solving for the optimal learning rate $\eta_{t}$. However, notice that the reverse is not true: any given schedule $\eta_{t}$ can be represented by a number of different weights $w_{t}$, and these weights give rise to different bounds using Theorem 1. The proof of Theorem 9 works by constructing a particular set of weights $w_{t}$, but these may not be the best weights in terms of providing the tightest convergence bound for the given $\eta_{t}$.
Although Theorem 9 shows that the learning rate representation of Theorem 3 indeed covers all possible schedules, it does not provide a user-friendly way to analyze the convergence of an arbitrary schedule. In this section, we provide an alternative analysis that fills this gap.
This approach is closer to previous final-iterate analysis techniques: we bound the final iterate in terms of the average iterate, plus an additional error term. This bound is strictly looser than those of Theorems 1 and 3 in the constant factors, but provides a convenient way to analyze arbitrary schedules. The bound is presented in Theorem 10 below.
Theorem 10
Suppose that $f$ is convex and let $x_{t}$ be given by SGD with learning rates $\eta_{t}$: $x_{t+1}=x_{t}-\eta_{t}g_{t}$. Then for the last iterate $x_{T}$ we have:
$\mathbb{E}[f(x_{T})-f(u)]$ $\leq\mathbb{E}\left[\frac{1}{2\sum_{t=1}^{T}\eta_{t}}D^{2}+\frac{1}{2\sum_{t=1}^{T}\eta_{t}}\sum_{t=1}^{T}\eta_{t}^{2}\left\|g_{t}\right\|^{2}\right]$
$+\mathbb{E}\left[\frac{1}{2}\sum_{k=1}^{T-1}\frac{\eta_{k}}{\sum_{t=k+1}^{T}\eta_{t}}\left(\frac{1}{\sum_{t=k}^{T}\eta_{t}}\sum_{t=k}^{T}\eta_{t}^{2}\left\|g_{t}\right\|^{2}\right)\right].$ (5)
Optimizing this bound with respect to the step-size sequence produces schedules that are visually indistinguishable from those of the regret based approach described in Section 2.1 for large $T$.
To prove Theorem 10, we will need to control quantities like:
$\sum_{t=k}^{T}\eta_{t}\left(q_{t}-q_{k}\right)=f(x_{t})-f(x_{k}).$
This can be bounded by the usual suboptimality inequality for SGD/GD methods, which holds for any $u$:
$\sum_{t=k}^{T}\eta_{t}\left[f(x_{t})-f(u)\right]\leq\frac{1}{2}\left\|x_{k}-u\right\|^{2}+\frac{1}{2}\sum_{t=k}^{T}\eta_{t}^{2}\left\|g_{t}\right\|^{2},$
Setting $u=x_{k}$ yields:
$\sum_{t=k}^{T}\eta_{t}\left[q_{t}-q_{k}\right]\leq\frac{1}{2}\sum_{t=k}^{T}\eta_{t}^{2}\left\|g_{t}\right\|^{2}.$
We can use this in our Lemma 5 to obtain the following result:
Corollary 11
$q_{T}=\frac{1}{\sum_{t=1}^{T}\eta_{t}}\sum_{t=1}^{T}\eta_{t}q_{t}+\frac{1}{2}\sum_{k=1}^{T-1}\frac{\eta_{k}}{\sum_{t=k+1}^{T}\eta_{t}}\left(\frac{1}{\sum_{t=k}^{T}\eta_{t}}\sum_{t=k}^{T}\eta_{t}^{2}\left\|g_{t}\right\|^{2}\right).$
Now, we are ready to prove Theorem 10. See 10
Proof  Following the standard convergence bound approach:
$\left\|x_{t+1}-u\right\|^{2}$ $=\left\|x_{t}-\eta_{t}g_{t}-u\right\|^{2}$
$=\left\|x_{t}-u\right\|^{2}-2\eta_{t}\left\langle g_{t},x_{t}-u\right\rangle+\eta_{t}^{2}\left\|g_{t}\right\|^{2}$
$\leq\left\|x_{t}-u\right\|^{2}-2\eta_{t}\left[f(x_{t})-f(u)\right]+\eta_{t}^{2}\left\|g_{t}\right\|^{2}.$
Summing over $t$ and telescoping gives:
$\sum_{t=1}^{T}\eta_{t}\left[f(x_{t})-f(u)\right]\leq\frac{1}{2}D^{2}+\sum_{t=1}^{T}\eta_{t}^{2}\left\|g_{t}\right\|^{2}.$
Then divide through by $\sum_{t=1}^{T}\eta_{t}$:
$\frac{1}{\sum_{t=1}^{T}\eta_{t}}\sum_{t=1}^{T}\eta_{t}\left[f(x_{t})-f_{*}\right]\leq\frac{1}{2\sum_{t=1}^{T}\eta_{t}}D^{2}+\frac{1}{2\sum_{t=1}^{T}\eta_{t}}\sum_{t=1}^{T}\eta_{t}^{2}\left\|g_{t}\right\|^{2}.$
Now we apply Corollary 11 to get:
$f(x_{T})-f_{*}$ $\leq\frac{1}{2\sum_{t=1}^{T}\eta_{t}}D^{2}+\frac{1}{2\sum_{t=1}^{T}\eta_{t}}\sum_{t=1}^{T}\eta_{t}^{2}\left\|g_{t}\right\|^{2}$
$+\frac{1}{2}\sum_{k=1}^{T-1}\frac{\eta_{k}}{\sum_{t=k+1}^{T}\eta_{t}}\left(\frac{1}{\sum_{t=k}^{T}\eta_{t}}\sum_{t=k}^{T}\eta_{t}^{2}\left\|g_{t}\right\|^{2}\right).$
By using a worst-case bound of $\|g_{t}\|^{2}\leq G^{2}$, we obtain:
Corollary 12
If $f$ is $G$-Lipschitz, then
$f(x_{T})-f_{*}$ $\leq\frac{1}{2\sum_{t=1}^{T}\eta_{t}}D^{2}+\frac{G^{2}}{2\sum_{t=1}^{T}\eta_{t}}\sum_{t=1}^{T}\eta_{t}^{2}$
$\quad+\frac{G^{2}}{2}\sum_{k=1}^{T-1}\frac{\eta_{k}}{\sum_{t=k+1}^{T}\eta_{t}}\left(\frac{1}{\sum_{t=k}^{T}\eta_{t}}\sum_{t=k}^{T}\eta_{t}^{2}\right).$
F.1 Linear schedule analysis with Theorem 10
In this section, we use Theorem 10 to re-analyze the the linear decay schedule:
$\eta_{t}=\frac{D}{G\sqrt{T}}\left(1-\frac{t}{T+1}\right).$ (6)
The resulting convergence rate is asymptotically correct, but does not achieve the optimal constant obtained by Theorem 3.
Theorem 13
Schedule 6 gives the following bound on the last iterate:
$f(x_{T})-f_{*}\leq\left(2+\frac{1}{4}\right)\frac{DG}{\sqrt{T}},$
or more precisely:
$f(x_{T})-f_{*}\leq\left(2+\frac{H\left(T-1\right)-2/3}{T+1}\right)\frac{DG}{\sqrt{T}},$
where H(T) is the Tth harmonic sum.
Proof  We start with the above bound:
$f(x_{T})-f_{*}$ $\leq\frac{1}{2\sum_{t=1}^{T}\eta_{t}}D^{2}+\frac{G^{2}}{2\sum_{t=1}^{T}\eta_{t}}\sum_{t=1}^{T}\eta_{t}^{2}$
$+\frac{G^{2}}{2}\sum_{k=1}^{T-1}\frac{\eta_{k}}{\sum_{t=k+1}^{T}\eta_{t}}\left(\frac{1}{\sum_{t=k}^{T}\eta_{t}}\sum_{t=k}^{T}\eta_{t}^{2}\right).$
Since the $\frac{D}{G\sqrt{T}}$ part of the step size is constant, it can be pulled outside the summations, so we just need to focus on summations involving $\left(1-\frac{t}{T+1}\right)$. For the first two terms we have in the bound, they simplify
$\sum_{t=1}^{T}\eta_{t}$ $\propto\sum_{t=1}^{T}\left(1-\frac{t}{T+1}\right)$
$=\left(T-\frac{1}{T+1}\sum_{t=1}^{T}t\right)$
$=\left(T-\frac{1}{2}\frac{T(T+1)}{T+1}\right)$
$=\frac{T}{2}.$
Similarly,
$\sum_{t=1}^{T}\eta_{t}^{2}$ $\propto\sum_{t=1}^{T}\left(1-\frac{t}{T+1}\right)^{2}$
$=\sum_{t=1}^{T}\left(1-\frac{2t}{T+1}+\frac{t^{2}}{T+1^{2}}\right)$
$=\left(T-T+\sum_{t=1}^{T}\frac{t^{2}}{T^{2}}\right)$
$=\left(\frac{T(T+1)(2T+1)}{6(T+1)^{2}}\right)$
$=\left(\frac{T(T+1)(T+1/2)}{3(T+1)^{2}}\right)$
$\leq\frac{T}{3}.$
So we have the following bound on the last iterate:
$f(x_{T})-f_{*}$ $\leq\frac{DG}{\sqrt{T}}+\frac{DG}{3\sqrt{T}}$
$+\frac{G^{2}}{2}\sum_{k=1}^{T-1}\frac{\eta_{k}}{\sum_{t=k+1}^{T}\eta_{t}}\left(\frac{1}{\sum_{t=k}^{T}\eta_{t}}\sum_{t=k+1}^{T}\eta_{t}^{2}\right).$
To simplify the remaining term we rely on computer algebra software. SymPy gives:
$\sum_{k=1}^{T-1}\frac{\eta_{k}}{\sum_{t=k+1}^{T}\eta_{t}}\left(\frac{1}{\sum_{t=k}^{T}\eta_{t}}\sum_{t=k+1}^{T}\eta_{t}^{2}\right)=\left(\frac{D}{G\sqrt{T}}\right)\frac{4T+6H\left(T-1\right)-4}{3(T+1)}.$
Where $H(T)=\sum_{t=1}^{T}1/t$ is the harmonic sum. So:
$f(x_{T})-f_{*}$ $\leq\frac{DG}{\sqrt{T}}+\frac{DG}{3\sqrt{T}}+\frac{2DG\sqrt{T}}{3T}+\frac{DG\left(3H\left(T-1\right)-2\right)}{3\left(T+1\right)\sqrt{T}}$
$\leq\frac{2DG}{\sqrt{T}}+\frac{DG}{3\sqrt{T}}+\frac{DG\left(3H\left(T-1\right)-2\right)}{3(T+1)\sqrt{T}}.$
Note that the term $\left(3H\left(T-1\right)-2\right)/3\left(T+1\right)\leq 1/4$ for all $T$, so:
$\frac{DG\left(3H\left(T-1\right)-2\right)}{3(T+1)\sqrt{T}}$ $=\frac{DG}{4\sqrt{T}},$
So combining with the $\frac{DG}{\sqrt{T}}+\frac{DG}{3\sqrt{T}}$ terms, we have:
$f(x_{T})-f_{*}\leq\left(2+\frac{1}{4}\right)\frac{DG}{\sqrt{T}}.$
bounding the harmonic function with a log gives instead:
$f(x_{T})-f_{*}\leq\frac{2DG}{\sqrt{T}}+O\left(\frac{DG\log(T)}{T^{3/2}}\right).$
Appendix G Experimental Setup
Our experiments on CIFAR-10, CIFAR-100, ImageNet and RCNN use SGD, and the remaining problems use Adam. We used decoupled weight decay with Adam in each case following standard practice for each problem.
G.1 Convex Experiments
Each dataset is obtained from the LIBSVM repository with used without modifications.
Hyper-parameter Value
GPUs 1$\times$×V100
Batch size 16
Epochs 100
Seeds 10
Hyper-parameter Value
Decay 0.0
Optimizer Adam
$\beta_{1}$ 0.9
$\beta_{2}$ 0.95
Figure 6: Logistic regression schedules, generated using a linear decay schedule with warmup for the initial run.
G.2 CIFAR-10
We used custom training code based on the PyTorch tutorial code for this problem. Following standard data-augmentation practises, we appliyed random horizontal flips and random offset cropping down to 32x32, using reflection padding of 4 pixels. Input pixel dataD was normalized by centering around 0.5.
Hyper-parameter Value
Architecture Wide ResNet 16-8
Epochs 300
GPUs 1$\times$×V100
Batch size per GPU 128
Hyper-parameter Value
Seeds 10
decay 0.0001
Momentum 0.9
G.3 CIFAR-100
We used the same codebase as for our CIFAR-10 experiments, with the same data augmentation.
We normalized each input image using fixed mean and standard error values derived from pre-processing the data.
Hyper-parameter Value
Architecture
DenseNet [6,12,24,16],
growth rate 12
Epochs 300
GPUs 1$\times$×V100
Hyper-parameter Value
Batch size per GPU 64
Seeds 10
Decay 0.0002
Momentum 0.9
G.4 ImageNet
We used the same code-base as for our CIFAR-10 experiments, and applied the same preprocessing procedure. The data-augmentations consisted of PyTorch’s RandomResizedCrop, cropping to 224x224 followed by random horizontal flips. Test images used a fixed resize to 256x256 followed by a center crop to 224x224.
Hyper-parameter Value
Architecture ResNet50
Epochs 100
GPUs 8$\times$×V100
Batch size per GPU 32
Hyper-parameter Value
Seeds 5
Decay 0.0001
Momentum 0.9
G.5 IWSLT14
We used the FairSeq framework  [^1] https://github.com/facebookresearch/fairseq for our experiments here, as well as for our GPT and RoBERTa experiments. Rather than a vanilla LSTM we use the variant from (Wiseman and Rush, 2016) provided in the FairSeq codebase.
Hyper-parameter Value
Architecture lstm_wiseman_iwslt_de_en
Max Epoch 55
GPUs 1$\times$×V100
Tokens per batch 4096
Warmup steps 4000
Dropout 0.3
Label smoothing 0.1
Hyper-parameter Value
Share decoder, input,
output embed
True
Float16 True
Update Frequency 1
Seeds 10
Decay 0.05
$\beta_{1},\beta_{2}$ 0.9, 0.98
G.6 RoBERTa
The RoBERTa implementation in FairSeq is the canonical one. We differ from the paper’s results by training for a shorter duration, which is necessary to keep our experiments computationally tractable. Our BookWiki dataset matches the original paper.
Hyper-parameter Value
Architecture roberta_base
Task masked_lm
Max updates 23,000
GPUs 8$\times$×V100
Max tokens per sample 512
Dropout 0.1
Attention Dropout 0.1
Max sentences 16
Hyper-parameter Value
Warmup 10,000
Sample Break Mode Complete
Float16 True
Update Frequency 16
Seeds 5
Decay 0.0
$\beta_{1},\beta_{2}$ 0.9, 0.98
G.7 GPT
Since the training dataset for GPT models are not availiable, we use the BookWiki dataset as used for RoBERTa training. Our model here is small, using 12 decoding layers and a decoder embedding dim of 768, giving 162 million parameters.
Hyper-parameter Value
Architecture transformer_lm_gpt
Task language_modeling
Max updates 65,000
GPUs 8$\times$×V100
Tokens per sample 512
Dropout 0.1
Attention Dropout 0.1
Max sentences 1
Warmup 10,000
Hyper-parameter Value
Sample Break Mode Complete
Share decoder, input,
output embed
True
Float16 True
Update Frequency 16
Seeds 5
Decay 0.005
$\beta_{1},\beta_{2}$ 0.9, 0.98
G.8 ViT
Our implementation uses the PyTorch Image Models library  [^2] https://github.com/rwightman/pytorch-image-models, with hyper-parameters following examples given in the repository.
Hyper-parameter Value
Model vit_tiny_patch16_224
GPUs 8$\times$×V100
Epochs 300
Batch Size 512
Warmup epochs 5
Hflip 0.5
aa rand-m6-mstd0.5
mixup 0.1
Hyper-parameter Value
mixup 0.1
cutmix 1.0
Crop Pct 0.9
BCE Loss True
Seeds 5
Decay 0.1
$\beta_{1},\beta_{2}$ 0.9, 0.999
G.9 DLRM
We used a custom implementation of the DLRM model based on the publicly available code. Our optimizer uses dense gradients for implementation simplicity, although sparse-gradients using AdaGrad is a more common baseline on this problem, we consider AdaGrad variants of our scheduling approach as future work.
Hyper-parameter Value
Iterations 300 000
Batch Size 128
Emb Dimension 16
GPUs 8$\times$×V100
Hyper-parameter Value
Seeds 5
Decay 0.0
$\beta_{1},\beta_{2}$ 0.9, 0.999
G.10 MRI
We used the version of the the fastMRI code base at https://github.com/facebookresearch/fastMRI/tree/main/banding_removal. Note that we found that training failed using PyTorch 2 or newer, and so we ran these experiments using PyTorch 1.9.
Hyper-parameter Value
Architecture 12 layer VarNet 2.0
Epochs 50
GPUs 8$\times$×V100
Batch size per GPU 1
Acceleration factor 4
Hyper-parameter Value
Low frequency lines 16
Mask type Offset-1
Seeds 5
Decay 0.0
$\beta_{1},\beta_{2}$ 0.9, 0.999
G.11 RCNN
Our RCNN experiments use Detectron2333https://github.com/facebookresearch/detectron2, and we use a pretrained ResNet backbone444 detectron2://ImageNetPretrained/MSRA/R-50.pkl.
Hyper-parameter Value
Backbone ResNet-50
Max Iter 200000
IMS Per Batch 16
Hyper-parameter Value
GPUs 8$\times$×V100
Momentum 0.9
Decay 1.5e-4